A glass tube (open at both ends) of length L is positioned near an audio speaker of frequency f = 680Hz. For what values of L will the tube resonate with the speaker?




1
Expert's answer
2021-02-22T16:04:48-0500

L is the length of the tube

f is the frequency

"\lambda" is the wavelength



"\dfrac{\lambda}{2}=d_{\lambda\lambda}=\dfrac{L}{n}"


"\Rightarrow L= \dfrac{n\lambda}{2}" for n= 1,2,3, K


But, we know that

"\lambda = \dfrac{v}{f}"


Putting the values of "\lambda," we get

"L=n(\dfrac{v}{2f})" for n= 1,2,3, K


"v=343" m/s and f= 680 Hz

On substituting, we get

"L =n(\dfrac{343m/s}{2\times 680Hz})=n(0.252)"


for n = 1,2,3,K


Hence, the value of L will be 0.252m , 0.504m, 0.756m ....